OpenStudy (anonymous):

checking my answer :) the region bounded by the parabola y=x^2 ,the line x=1 ,and the x-axis is rotated about the y-axis. Find the volume of the resulting solid

OpenStudy (anonymous):

V=\[\int\limits_{0}^{1}\Pi(x ^{2})^{2}dx \]

OpenStudy (anonymous):

=\[\pi(\frac{ x ^{5} }{ 5 })\]

OpenStudy (anonymous):

from 0 to 1

OpenStudy (anonymous):

I got pi/5

zepdrix (zepdrix):

Woops! If you're integrating in x, it looks like we are getting shells not disks, yes?